OpenStudy (anonymous):

im really bad with this slope intercept form :( write an equation in slope intercept form for the line that satisfies the following condition slope 7/18 and passes through (4,-20)

OpenStudy (ksaimouli):

y=mx+b

OpenStudy (ksaimouli):

me too even iam ad toooooo

OpenStudy (anonymous):

could u show me how to set it up

OpenStudy (anonymous):

\[m=\frac{y-y_1}{x-x_1}\] \[m=\frac7{18}\] \[y_1=-20\] \[x_1=4\] \[\frac7{18}=\frac{y-(-20)}{x-4}\] you can solve the rest right?

OpenStudy (shayaan_mustafa):

as slope-point formula is given as, m=(y-y1)/(x-x1) m=7/18 x1=4 y1=-20 then, 7/18 = (y-4)/(x+20) 7x + 140 = 18y - 72 18y = -7x -72-140 18y = -7x -212 making its standard form as, y=mx+b y = (-7/18)x - (212) Regards, Read my Profile. :)

OpenStudy (anonymous):

\[y - y1 = m \times (x - x1)\] Just insert m=7/18 x1 = 4 and y1 = -20